The people on thohoyandou island buy only juice and cloth the cpi basket contains the quantities bought in 2010 the average household spent $60 on juice and $30 on cloth in 2009 when the price of juice was $2 a bottle and the price of cloth was $5 a meter in the current year 2010 juice is $4 a bottle and cloth is $6 a meter calculate cpi per basket [3]
1
Expert's answer
2019-08-14T09:31:41-0400

2009= $60 on juice and $30 on cloth,juice was $2 a bottle and the price of cloth was $5 a meter = 30 bottles and 6 meters of cloth= COST OF GOODS= 90

2010= juice is $4 a bottle and cloth is $6 a meter= the average household spent= 4"\times" 30= $120 on juice and 6"\times"6=36 on cloth = COST OF GOODS= 156

CPI= "\frac{156}{90}" "\times" 100= 173.33